newton's second law


1
Expert's answer
2020-07-20T15:16:42-0400

According to the Second Newton's Law


"\sum \overrightarrow{F}=m\overrightarrow{a}=\frac{d\overrightarrow{p}}{dt}"


"\sum \overrightarrow{F}" - vector sum of all forces acting on the body


"m" - mass of the body


"\overrightarrow{a}" - acceleration of the body


"d\overrightarrow{p}" - change in momentum of the body during "dt"